# TAT-LLM: A Specialized Language Model for Discrete Reasoning over Tabular and Textual Data

### v1

@@ 4.6. Model Analysis
Extractor Wrong Evidence (48%)
Q: What is the percentage change in cash flow hedges in 2011 compared to 2010?
@@
Missed Evidence (15%)
Q: What was the total impairment costs recorded from 2003 to 2005 in millions?
@@
Redundant Evidence (8%)
Q: What was the average number of shares issued to employees from 2013 to 2015?
@@
Reasoner Wrong Operators (19%)
Q: what was the percent of the change in the stock price performance for hum from 2010 to 2011?
@@
Wrong Values (10%)
Q: What was the average cash flow from 2004 to 2006?
@@
4.6.5. Error Analysis
To further diagnose our TAT-LLM model, we randomly sample $100$ error instances of TAT-LLM (7B) from the test set of FinQA and analyze the reasons. Since we adopt the External Executor for reliable execution, the errors only occur to the Extractor and Reasoner. As shown in Table 7, we list all kinds of errors with examples and their corresponding percentages. We can observe that most of the errors come from the Extractor, where the Wrong Evidence takes $48\%$ of the total errors. This shows that our TAT-LLM (7B) model still faces challenges in precisely interpreting the meaning of the values, sometimes due to the unique terminology. Training the model with more tabular and textual data in this domain might be a possible approach to enhancing its evidence extraction ability. Additionally, the Wrong Operator and the Wrong Values may be partially caused by the randomness of the LLM as the generation length increases, which fails to generate the full equation or wrongly copies the extracted evidence.
5. Conclusion
In this work, we first abstract a Step-wise Pipeline for tabular and textual QA based on the previous multi-step methods. Following this pipeline, we develop the TAT-LLM model by specializing smaller language models (i.e., LLaMA 2) in discrete reasoning over tabular and textual data. We validate its effectiveness with extensive experiments and analyses, which show that our TAT-LLM model can outperform both conventional methods and very large-scale LMs like GPT-4 on this task. Our work well demonstrates that specialized models might be a promising direction towards more advanced models on specific tasks that can compete with human expertise.
Appendix A Templates for Fine-tuning
@@
Table 1. The template for constructing training instances following Step-wise Pipeline with FinQA dataset.
@@
### Instruction:
Given a table and a list of texts in the following, what is the answer to the question? Please complete the task in three steps:
1. In the first step, extract the relevant numerical values from the provided table or texts. Store these in the variable ‘{evidence}‘. If there are multiple values, separate them using the ’#’ symbol.
2. In the second step, generate an equation using the extracted numerical values. Store this equation in the variable ‘{equation}‘.
3. In the third step, calculate the answer based on the equation and store it in the variable ‘{answer}‘.
Please organize the results in the following table:
— step — output —
— 1 — {evidence} —
— 2 — {equation} —
— 3 — {answer} —
Finally, present the calculated answer in the format: ”The answer is: {answer}”
### Table
{table}
### Text
{text}
### Question
{question}
### Response
—step — output—
—1 — {gold_evidence} —
—2 — {gold_equation} —
—3 — {gold_answer} —
@@
Table 2. The template for constructing training instances following Step-wise Pipeline with TAT-QA dataset.
@@
Given a table and a list of texts in the following, answer the question posed using the following five-step process:
1. Step 1: Predict the type of question being asked. Store this prediction in the variable ‘{question_type}‘. The value of ‘{question_type}‘ can be one of the following:‘Single span‘, ‘Multiple spans‘, ‘Count‘, or ‘Arithmetic‘.
2. Step 2: Extract the relevant strings or numerical values from the provided table or texts. Store these pieces of evidence in the variable ‘{evidence}‘. If there are multiple pieces of evidence, separate them using the ’#’ symbol.
3. Step 3: if the ‘{question_type}‘ is ‘Arithmetic‘, formulate an equation using values stored in ‘{evidence}‘. Store this equation in the variable ‘{equation}‘. For all other question types, set the value of {equation} to ’N.A.’.
4. Step 4: Predict or calculate the answer based on the question type, evidence and equation. Store it in the variable ‘{answer}‘. If there are multiple values, separate them using the ’#’ symbol.
5. Step 5: If the value of the ‘{answer}‘ is numerical, predict its scale and store it in a variable named ‘{scale}‘. The value of ‘{scale}‘ can be one of the following: ‘none‘, ‘percent‘, ‘thousand‘, ‘million‘, or ‘billion‘. For non-numerical values, set the value of ‘{scale}‘ to ’none’.
Please organize the results in the following table:
— step — output —
— 1 — {question_type} —
— 2 — {evidence} —
— 3 — {equation} —
— 4 — {answer} —
— 5 — {scale} —
Finally, present the final answer in the format: ”The answer is: {answer} #### and its corresponding scale is: {scale}”
### Table
{table}
### Text
{text}
### Question
{question}
### Response
— step — output —
— 1 — {gold_question_type} —
— 2 — {gold_evidence} —
— 3 — {gold_equation} —
— 4 — {gold_answer} —
— 5 — {gold_scale} —
Table 3. The template for constructing training instances following Step-wise Pipeline with TAT-DQA dataset.
### Instruction
Given a document that has one or multiple pages in the following, answer the question posed using the following five-step process:
1. Step 1: Predict the type of question being asked. Store this prediction in the variable ‘{question_type}‘. The value of ‘{question_type}‘ can be one of the following:‘Single span‘, ‘Multiple spans‘, ‘Count‘, or ‘Arithmetic‘.
2. Step 2: Extract the relevant strings or numerical values from the provided document. Store these pieces of evidence in the variable ‘{evidence}‘. If there are multiple pieces of evidence, separate them using the ’#’ symbol.
3. Step 3: if the ‘{question_type}‘ is ‘Arithmetic‘, formulate an equation using values stored in ‘{evidence}‘. Store this equation in the variable ‘{equation}‘. For all other question types, set the value of {equation} to ’N.A.’.
4. Step 4: Predict or calculate the answer based on the question type, evidence and equation. Store it in the variable ‘{answer}‘. If there are multiple values, separate them using the ’#’ symbol.
5. Step 5: If the value of the ‘{answer}‘ is numerical, predict its scale and store it in a variable named ‘{scale}‘. The value of ‘{scale}‘ can be one of the following: ‘none‘, ‘percent‘, ‘thousand‘, ‘million‘, or ‘billion‘. For non-numerical values, set the value of ‘{scale}‘ to ’none’.
Please organize the results in the following table:
— step — output —
— 1 — {question_type} —
— 2 — {evidence} —
— 3 — {equation} —
— 4 — {answer} —
— 5 — {scale} —
Finally, present the final answer in the format: ”The answer is: {answer} #### and its corresponding scale is: {scale}”
### Text
{pages}
### Question
{question}
### Response
— step — output —
— 1 — {gold_question_type} —
— 2 — {gold_evidence} —
— 3 — {gold_equation} —
— 4 — {gold_answer} —
— 5 — {gold_scale} —
Table 4. The template for constructing training instances following End-to-end Pipeline with FinQA dataset
### Instruction
Given a table and a list of texts in the following, what is the answer to the question? Please output the answer in the format of ”The answer is:”.
@@
Table 5. The template for constructing training instances following End-to-end Pipeline with TAT-QA dataset
@@
Given a table and a list of texts in the following, what is the answer to the question? Please predict the answer and store it in a variable named ‘{answer}‘. If there are multiple values, separate them using the ’#’ symbol. If the value of the ‘{answer}‘ is numerical, predict its scale and store it in a variable named ‘{scale}‘. The value of ‘{scale}‘ can be one of the following: ‘none‘, ‘percent‘, ‘thousand‘, ‘million‘, or ‘billion‘. For non-numerical values, set the value of ‘{scale}‘ to ’none’. Finally, present the final answer in the format of ”The answer is: {answer} #### and its corresponding scale is: {scale}”
### Table
{table}
### Text
{text}
@@
### Response
Table 6. The template for constructing training instances following End-to-end Pipeline with TAT-DQA dataset
@@
Given a document that has one or multiple pages in the following, what is the answer to the question? Please predict the answer and store it in a variable named ‘{answer}‘. If there are multiple values, separate them using the ’#’ symbol. If the value of the ‘{answer}‘ is numerical, predict its scale and store it in a variable named ‘{scale}‘. The value of ‘{scale}‘ can be one of the following: ‘none‘, ‘percent‘, ‘thousand‘, ‘million‘, or ‘billion‘. For non-numerical values, set the value of ‘{scale}‘ to ’none’. Finally, present the final answer in the format of ”The answer is: {answer} #### and its corresponding scale is: {scale}”
### Document
{pages}
@@
Appendix B Templates for Zero-shot Inference
For zero-shot inference with the baseline LLMs such as GPT-3.5-Turbo (Brown et al., 2020) and GPT-4 (OpenAI, 2023), we utilize three different templates to build the instances and feed them to the LLMs to get the predictions. We report the best results in Section 4.5. The following are details of the templates we used. 1) End-to-end Pipeline template, i.e., Table A7 for FinQA and TAT-QA, and Table A8 for TAT-DQA; 2) Chain-of-Thought template, which adds ”Let’s complete this task step by step” to the End-to-end Pipeline template; 3) Step-wise Pipeline template that is similar to the templates for fine-tuning, with the part of “Response” removed.
Table 7. The template for zero-shot inference using the baseline LLMs on FinQA and TAT-QA datasets.
### Instruction
Given a table and a list of texts in the following, what is the answer to the question? (Let’s complete this task step by step.) Please output the answer in the format of ”The answer is:”.
### Table
{table}
### Text
{text}
### Question
{question}
@@ Appendix B Templates for Zero-shot Inference
Table 8. The template for zero-shot inference using the baseline LLMs on TAT-DQA datasets.
### Instruction
Given a document with one or multiple pages in the following, what is the answer to the question? (Let’s complete this task step by step.) Please output the answer in the format of ”The answer is:”.
### Text
{pages}
### Question
{question}
### Response

### v3

@@ Appendix A Implementation Details
Extractor Wrong Evidence (48%) Q: What is the percentage change in cash flow hedges in 2011 compared to 2010?
@@
Missed Evidence (15%) Q: What was the total impairment costs recorded from 2003 to 2005 in millions?
@@
Redundant Evidence (8%) Q: What was the average number of shares issued to employees from 2013 to 2015?
@@
Reasoner Wrong Operators (19%) Q: what was the percent of the change in the stock price performance for hum from 2010 to 2011?
@@
Wrong Values (10%) Q: What was the average cash flow from 2004 to 2006?
@@
Table 1: Examples of errors on the test set of FinQA and corresponding percentages in each step. Q, G and P denote question, ground truth and prediction.
Appendix B Error Analysis
To further diagnose our TAT-LLM model, we randomly sample $100$ error instances of TAT-LLM (7B) from the test set of FinQA and analyze the reasons. Since we adopt the External Executor for reliable execution, the errors only occur to the Extractor and Reasoner. As shown in Table A1, we list all kinds of errors with examples and their corresponding percentages. We can observe that most of the errors come from the Extractor, where the Wrong Evidence takes $48\%$ of the total errors. This shows that our TAT-LLM (7B) model still faces challenges in precisely interpreting the meaning of the values, sometimes due to the unique terminology. Training the model with more tabular and textual data in this domain might be a possible approach to enhancing its evidence extraction ability. Additionally, the Wrong Operator and the Wrong Values may be partially caused by the randomness of the LLM as the generation length increases, which fails to generate the full equation or wrongly copies the extracted evidence.
Appendix C Templates for Fine-tuning
@@
### Instruction: Given a table and a list of texts in the following, what is the answer to the question? Please complete the task in three steps: 1. In the first step, extract the relevant numerical values from the provided table or texts. Store these in the variable ‘{evidence}‘. If there are multiple values, separate them using the ’#’ symbol. 2. In the second step, generate an equation using the extracted numerical values. Store this equation in the variable ‘{equation}‘. 3. In the third step, calculate the answer based on the equation and store it in the variable ‘{answer}‘.
Please organize the results in the following table: | step | output | | 1 | {evidence} | | 2 | {equation} | | 3 | {answer} |
Finally, present the calculated answer in the format: "The answer is: {answer}"
### Table {table}
### Text {text}
### Question {question}
### Response |step | output| |1 | {gold_evidence} | |2 | {gold_equation} | |3 | {gold_answer} |
@@
Table 2: The template for constructing training instances following Step-wise Pipeline with FinQA dataset.
### Instruction Given a table and a list of texts in the following, answer the question posed using the following five-step process:
1. Step 1: Predict the type of question being asked. Store this prediction in the variable ‘{question_type}‘. The value of ‘{question_type}‘ can be one of the following:‘Single span‘, ‘Multiple spans‘, ‘Count‘, or ‘Arithmetic‘. 2. Step 2: Extract the relevant strings or numerical values from the provided table or texts. Store these pieces of evidence in the variable ‘{evidence}‘. If there are multiple pieces of evidence, separate them using the ’#’ symbol. 3. Step 3: if the ‘{question_type}‘ is ‘Arithmetic‘, formulate an equation using values stored in ‘{evidence}‘. Store this equation in the variable ‘{equation}‘. For all other question types, set the value of {equation} to ’N.A.’. 4. Step 4: Predict or calculate the answer based on the question type, evidence and equation. Store it in the variable ‘{answer}‘. If there are multiple values, separate them using the ’#’ symbol. 5. Step 5: If the value of the ‘{answer}‘ is numerical, predict its scale and store it in a variable named ‘{scale}‘. The value of ‘{scale}‘ can be one of the following: ‘none‘, ‘percent‘, ‘thousand‘, ‘million‘, or ‘billion‘. For non-numerical values, set the value of ‘{scale}‘ to ’none’.
Please organize the results in the following table: | step | output | | 1 | {question_type} | | 2 | {evidence} | | 3 | {equation} | | 4 | {answer} | | 5 | {scale} |
Finally, present the final answer in the format: "The answer is: {answer} #### and its corresponding scale is: {scale}"
### Table {table}
### Text {text}
### Question {question}
### Response | step | output | | 1 | {gold_question_type} | | 2 | {gold_evidence} | | 3 | {gold_equation} | | 4 | {gold_answer} | | 5 | {gold_scale} |
Table 3: The template for constructing training instances following Step-wise Pipeline with TAT-QA dataset.
### Instruction Given a document that has one or multiple pages in the following, answer the question posed using the following five-step process:
1. Step 1: Predict the type of question being asked. Store this prediction in the variable ‘{question_type}‘. The value of ‘{question_type}‘ can be one of the following:‘Single span‘, ‘Multiple spans‘, ‘Count‘, or ‘Arithmetic‘. 2. Step 2: Extract the relevant strings or numerical values from the provided document. Store these pieces of evidence in the variable ‘{evidence}‘. If there are multiple pieces of evidence, separate them using the ’#’ symbol. 3. Step 3: if the ‘{question_type}‘ is ‘Arithmetic‘, formulate an equation using values stored in ‘{evidence}‘. Store this equation in the variable ‘{equation}‘. For all other question types, set the value of {equation} to ’N.A.’. 4. Step 4: Predict or calculate the answer based on the question type, evidence and equation. Store it in the variable ‘{answer}‘. If there are multiple values, separate them using the ’#’ symbol. 5. Step 5: If the value of the ‘{answer}‘ is numerical, predict its scale and store it in a variable named ‘{scale}‘. The value of ‘{scale}‘ can be one of the following: ‘none‘, ‘percent‘, ‘thousand‘, ‘million‘, or ‘billion‘. For non-numerical values, set the value of ‘{scale}‘ to ’none’.
Please organize the results in the following table: | step | output | | 1 | {question_type} | | 2 | {evidence} | | 3 | {equation} | | 4 | {answer} | | 5 | {scale} |
Finally, present the final answer in the format: "The answer is: {answer} #### and its corresponding scale is: {scale}"
### Text {pages}
### Question {question}
### Response | step | output | | 1 | {gold_question_type} | | 2 | {gold_evidence} | | 3 | {gold_equation} | | 4 | {gold_answer} | | 5 | {gold_scale} |
Table 4: The template for constructing training instances following Step-wise Pipeline with TAT-DQA dataset.
@@
Given a table and a list of texts in the following, what is the answer to the question? Please output the answer in the format of "The answer is:".
@@
Table 5: The template for constructing training instances following End-to-end Pipeline with FinQA dataset
@@
Given a table and a list of texts in the following, what is the answer to the question? Please predict the answer and store it in a variable named ‘{answer}‘. If there are multiple values, separate them using the ’#’ symbol. If the value of the ‘{answer}‘ is numerical, predict its scale and store it in a variable named ‘{scale}‘. The value of ‘{scale}‘ can be one of the following: ‘none‘, ‘percent‘, ‘thousand‘, ‘million‘, or ‘billion‘. For non-numerical values, set the value of ‘{scale}‘ to ’none’. Finally, present the final answer in the format of "The answer is: {answer} #### and its corresponding scale is: {scale}"
### Table {table}
### Text {text}
@@
### Response The answer is: {gold_answer} #### and its corresponding scale is: {gold_scale}
Table 6: The template for constructing training instances following End-to-end Pipeline with TAT-QA dataset
@@
Given a document that has one or multiple pages in the following, what is the answer to the question? Please predict the answer and store it in a variable named ‘{answer}‘. If there are multiple values, separate them using the ’#’ symbol. If the value of the ‘{answer}‘ is numerical, predict its scale and store it in a variable named ‘{scale}‘. The value of ‘{scale}‘ can be one of the following: ‘none‘, ‘percent‘, ‘thousand‘, ‘million‘, or ‘billion‘. For non-numerical values, set the value of ‘{scale}‘ to ’none’. Finally, present the final answer in the format of "The answer is: {answer} #### and its corresponding scale is: {scale}"
### Document {pages}
@@
### Response The answer is: {gold_answer} #### and its corresponding scale is: {gold_scale}
Table 7: The template for constructing training instances following End-to-end Pipeline with TAT-DQA dataset
Appendix D Templates for Zero-shot Inference
For zero-shot inference with the baseline LLMs such as GPT-3.5-Turbo Ouyang et al. (2022) and GPT-4 OpenAI (2023), we utilize three different templates to build the instances and feed them to the LLMs to get the predictions. We report the best results in Section 4.2. The following are details of the templates we used. 1) End-to-end Pipeline template, i.e., Table A8 for FinQA and TAT-QA, and Table A9 for TAT-DQA; 2) Chain-of-Thought template, which adds "Let’s complete this task step by step" to the End-to-end Pipeline template; 3) Step-wise Pipeline template that is similar to the templates for fine-tuning, with the part of “Response” removed.
### Instruction Given a table and a list of texts in the following, what is the answer to the question? (Let’s complete this task step by step.) Please output the answer in the format of "The answer is:".
### Table {table}
### Text {text}
### Question {question}
@@ Appendix D Templates for Zero-shot Inference
Table 8: The template for zero-shot inference using the baseline LLMs on FinQA and TAT-QA datasets.
### Instruction Given a document with one or multiple pages in the following, what is the answer to the question? (Let’s complete this task step by step.) Please output the answer in the format of "The answer is:".
### Text {pages}
### Question {question}
@@
Table 9: The template for zero-shot inference using the baseline LLMs on TAT-DQA datasets.